OpenStudy (anonymous):

f is nonnegative on the indicated interval, find the area of the region bounded by the graph of f and the x-axis. f(x)= (5x)/ (2x^2+1)^2 x E[0, 2]

OpenStudy (anonymous):

guess you have to integrate right?

OpenStudy (anonymous):

\[\int_0^2\frac{5x}{(2x^2+1)^2}dx\]

OpenStudy (anonymous):

a "u substitution" should work for this integral try \(u=2x^2+1,du=4xds, \frac{5}{4}du=5xdx\) while you are at it change the limits of integration \(u(0+1,u(2)=7\)

OpenStudy (anonymous):

i meant \(u(0)=1,u(2)=7\)

OpenStudy (anonymous):

you get \[\frac{5}{4}\int_1^7\frac{du}{u^2}\] and how you should be good to go

OpenStudy (anonymous):

Thanks! I just forgot to change the lim of i. Derp! :)